Can you help me with this equation?
If Ben invests $3400 at 5% interest per year, how much additional money must he invest at 6.5% annual interest to ensure that the interest he receives each year is 5.5% of the total amount invested?

Answer by checkley71(8403)

(3400+x).055=3400*.05+.065x
187+.055x=170+.065x
.055x-.065x=170-187
-.01x=-17
x=-17/-.01
x=$1,700 amount that needs to be invested @ 6.5%
proof
(3400+1700).055=3400*.05+.065*1700
5100*.055=170+110.5
280.5=280.5